Two dice are rolled repeatedly until the sum of the two numbers rolled is 10 or more.

a) What is the probability that exactly 5 rolls are needed? (Count each time you roll the dice as one roll).

b) What is the probability that more than 5 rolls are needed?

c) Find the expected number of rolls.
Answer #1

a)

here P(sum is 10 or more on pair of dice) =(P(10)+P(!1)+P(12)=(3/36)+(2/36)+(1/36) =6/36 =1/6

P( exactly 5 rolls are needed) =P(in first 4 rolls sum is less than 10 and on 5th 10 or more)

=(1-1/6)4*(1/6)=625/7776

b)

probability that more than 5 rolls are needed =P(first 5 roll all are below 10)=(5/6)6 =3125/7776

c)

expected number of rolls =1/p=1/(1/6)=6